OpenStudy (anonymous):

I need to find the equation of the line with a slope of 1/5 passing through the point (10,9). Any help?

sam (.sam.):

Use \(y-y_1=m(x-x_1)\), where 'm' is the slope, with coordinates \((x_1,y_1)\)

OpenStudy (anonymous):

Okay. So would the final equation be y=1/5x+7?

OpenStudy (anonymous):

@.Sam.

sam (.sam.):

Yup